OpenStudy (anonymous):

The inverse of the function y = x2 is still a function. a) Always b) Sometimes c) Never

myininaya (myininaya):

Well if you know point (4,16) and (-4,16) is on y=x^2 what two points from that do you know is on the inverse of y=x^2?

OpenStudy (anonymous):

no clue

myininaya (myininaya):

So you never heard of if (x,y) is on f, then (y,x) is on the inverse of f?

OpenStudy (anonymous):

i have not

myininaya (myininaya):

so what do you know about inverse functions then ?

myininaya (myininaya):

do you know how to find the inverse function algebraically?

OpenStudy (anonymous):

i know nothing about algebra

myininaya (myininaya):

That can't be true since this an algebra question.

OpenStudy (anonymous):

i copied it

myininaya (myininaya):

There isn't much I can do with nothing.

myininaya (myininaya):

You have to know something about inverse functions to answer this question. If you recall no definitions, then how can I help you.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

u could help by determining if it's a b or c

myininaya (myininaya):

oh no i can't just give the answer

myininaya (myininaya):

How about I try to teach you just a little about inverses?

myininaya (myininaya):

and functions.

myininaya (myininaya):

and then you can determine the answer

myininaya (myininaya):

First of all if you graph y=x^2 you have a parabola which it looks like  y=x^2 is a function I'm going to assume you know a little about algebra. I want you to reflect this graph about the y=x line and I will draw the y=x line so you know what line to reflect y=x^2 about.

myininaya (myininaya):